If the line ax+by+c=0 is a normal to the curve xy=1, then:
(A)a>0,b>0
(B)a>0,b<0
(C)a<0,b<0
(D)data is insufficient.

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Jayati

ax+by+c=0

slope of line m1 =-a/b

now for curve

xy=1

differentiate

xdy/dx +y=0

dy/dx =-y/x

 dy/dx =-1/x2

this is the slope of tangent

then slope of normal at any point is m =x2

so -a/b =x2

a and b should be apposite sign

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin